(1)  Suppose the First National Bank holds $900,000 in deposits and $300,000 as reserve:


(a)How much is left to be loaned out?

(b) What is the reserve ratio and multiplier of this transaction?



(2)  Suppose the First National Bank holds $700,000 in deposits and the multiplier is 25:


(a)Show the transactions of 1st, 2nd and 3rd National Bank.

(b)What is the total money supply in the economy?


1
Expert's answer
2021-01-12T13:04:47-0500

(1) If the First National Bank holds $900,000 in deposits and $300,000 as reserve:

(a) $900,000 is left to be loaned out?

(b) The reserve ratio is:

"300,000\/900,000 = 0.33."

And multiplier of this transaction is:

"m = 1\/0.33 = 3."

(2) If the First National Bank holds $700,000 in deposits and the multiplier is 25, and reserve ratio is 1/25 = 0.04:

(a) The transactions of 1st, 2nd and 3rd National Bank are:

First bank will put into reserves 700,000×0.04 = 28,000 and loan 672,000.

Second bank will put into reserves 672,000×0.04 = 26,880 and loan 645,120.

The third bank will put into reserves 645,120×0.04 = 25,804.8 and loan 619,315.2.

(b) The total money supply in the economy is 700,000×25 = 17,500,000.